Predict ridership for the next month if the Number of Tourists is expected to be 21 million? Use Excel. (2 points)

Month

Monthly Tourists (Forecast in Millions)

Monthly Uber Ridership (In Millions)

January

15

2.7

February

12

2.0

March

20

4.4

April

14

2.5

May

14

2.5

June

6

1.3

July

7

1.5

August

2

1.0

September

7

1.7

October

10

2.2

November

8

1.5

December

11

3.1
Answer #1

So we have number of tourists and based on this we need to predict Uber ridership. This means

Dependent variable (y values) = Monthly Uber ridership

Independent variable (x values) = Monthly tourists

We need to form an equation such as y = ax + b

Here

a = slope

b = intercept

On excel we can use formulas

=INTERCEPT(Y_range,X_range)

=SLOPE(Y_range,X_range)

The calculated values are shown below

The formulas are shown below

The equation is

Y = 0.3655 + 0.1747X

If the value of X = 21 then the ridership can be predicted as 0.3655 + 0.1747*21 = 4.03 million